### v1

@@ 1 Introduction
We repeated and improved upon the experiment conducted by the previous study [19], making targeted (yet as simple as possible) modifications to the original CRT problems. If LLMs genuinely possess the intrinsic capability to comprehend mathematical logic, as is widely hypothesized, their accuracy in responding to the modified queries should not experience a marked decrement. However, our experiments revealed a distinctly opposing result: even when employing the CoT approach, prominent LLMs, including the latest o1 model, continued to manifest a considerable error rate for the modified problems. Further analysis indicates that LLMs may have not developed a logic reasoning proficiency akin to System 2, nor have they acquired comprehensive mathematical cognitive abilities, including principles of mathematical operations and algorithm selection. They predominantly resort to a methodology reminiscent of System 1, which involves matching and producing responses based on the similarity between user inquiries and training data during the process of text generation. This investigation may serve to temper the overly optimistic anticipations regarding the effectiveness of CoT and the competencies of LLMs in approximating AGI.
@@ 2 Experiment I: Changing the numbers in a problem without altering its description and principle
We first replicated the experiment on the CRT3 dataset from Thilo Hagendorff’s study[19], which comprises 50 mathematical test problems. We introduced three types of modifications to the test problems, as delineated in Table 1. Type A represents the original problems in the CRT3 test, characterized as typical exponential growth problems, and incorporates three numbers in the problem statement. Type B modifies two of these numbers, while type C modifies all three. Type D replaces two numbers with letters, transforming the arithmetic problem into an algebraic one. To prevent LLMs from being unsure how to handle algebraic symbols, we appended a suffix for guidance: “X and Y are both numbers, you can use them to represent the final answer.” The modifications in Types B, C, and D do not alter the description of the original problem (Type A). Thus, the underlying mathematical principle of the original problem remains unchanged. To uncover the problem-solving strategies of LLMs, we appended explicit instructions (’Please think step by step’) to the end of each of the four types of problems, encouraging the use of the CoT method.
@@
We evaluated five mainstream open-source and commercial models, manually analyzing and recording their accuracy in responding to the problems. A response was considered correct only if both the problem-solving approach and the final answer were accurate (refer to Appendices Table 2 and Table 3 ). To ensure the reliability of our experimental results, we conducted three replicates of the experiments and computed the average for each data set. The results are shown in Figure 1. In the case of Type A problems, the highest accuracy reached 100.00%, the lowest was 54.00%, and the average accuracy was 86.80%. For Type B problems, which involved modifications to two numbers, the accuracy significantly decreased, with the highest being 92.00%, the lowest 25.33%, and the average 68.53%($\delta=18.27\%;\chi^{2}(1)=23.00;P<0.001$). For the problems of Type C, where three numbers were modified, the accuracy further declined, with the highest at 80.00%, the lowest at 10.67%, and the average 53.06%($\delta=15.47\%;\chi^{2}(1)=11.91;P<0.001$). The accuracy dropped sharply once again when the arithmetic problems were transformed into algebraic ones (Type D). The best-performing generic model, GPT-4, achieved an accuracy of only 29.33%, whereas ChatGPT 3.5 scarcely provided correct responses to any problems. The other models only got around ten problems right, reducing the average accuracy down to 20.93%($\delta=32.13\%;\chi^{2}(1)=54.00;P<0.001$).
@@ 4 Experiment III: Replicating Experiment I & II on the Latest o1 Model
However, when replicating Experiment II, the average accuracy rate of o1 was only 10%, which was even lower than that of most previous LLMs. After making minimal textual modifications while altering the mathematical principles underlying the original CRT1, CRT2, and CRT3 problems, o1 persisted in selecting problem-solving approaches based on the mathematical principles corresponding to the original problems. The errors stemming from this persistence constituted 100%, 68.42%, and 75.86% of the incorrect responses to the three types of CRT problems, respectively. This may indicate that technologies such as CoT, prompt engineering, and even fine-tuning tailored to specific datasets, cannot fundamentally enhance the ability of LLMs to comprehend mathematical problems. The reason for this lies in the fact that the learning paradigm of LLMs has not undergone substantial changes (such as those involving word chains), leading to their thought patterns being deeply ingrained.
@@ Appendix A Appendix
Problem Type Incorrect Answer Example
@@
A {”index”: 20, ”query”: ”A colony of bacteria is growing on yogurt, and the size of the colony doubles daily. If it takes 4 days for the colony to completely cover the yogurt, how many days does it take for the patch to cover half of the yogurt? Please think step by step.”, ”answer”: ”If the colony doubles in size every day, then on the first day it covers half of the yogurt. On the second day, it doubles in size and covers the other half of the yogurt, so now the yogurt is fully covered. Therefore, it takes 1 day for the colony to cover half of the yogurt.”, ”time”: ”2024-08-23 09:22:56”, ”Model”: ”gpt-3.5-turbo-0125”}
@@
B {”index”: 20, ”query”: ”A colony of bacteria is growing on yogurt, and the size of the colony triples daily. If it takes 4 days for the colony to completely cover the yogurt, how many days does it take for the patch to cover 1/27 of the yogurt? Please think step by step.”, ”answer”: ”Let’s break down the problem step by step: 1. The colony triples daily: This means that if the initial size of the colony is 1 unit, it will be 3 units the next day, 9 units the following day, 27 units the day after, and so on. 2. It takes 4 days for the colony to completely cover the yogurt: This means that the colony reaches its maximum size on the 4th day, which is 27 units. 3. We need to find out how many days it takes for the colony to cover 1/27 of the yogurt: Since the colony triples in size daily, on the third day, the colony would be 9 units, which is 1/3 of the complete coverage. So, it means that on the third day the yogurt is 1/3 covered. Therefore, it takes 3 days for the patch to cover 1/27 of the yogurt.”, ”time”: ”2024-08-13 13:22:38”, ”Model”: ”gpt-3.5-turbo-0125”}
@@
C {”index”: 20, ”query”: ”A colony of bacteria is growing on yogurt, and the size of the colony triples every three days. If it takes 27 days for the colony to completely cover the yogurt, how many days does it take for the patch to cover 1/27 of the yogurt? Please think step by step.”, ”answer”: ”First, let’s figure out how many times the colony triples in size in the 27 days. Since the colony triples every three days, it will triple in size 27/3 = 9 times. Now, we know that the colony starts at 1/27 of the yogurt and triples in size 9 times. So, to find out how many days it takes for the colony to cover 1/27 of the yogurt, we need to divide 27 by 9, which equals 3. Therefore, it takes 3 days for the colony to cover 1/27 of the yogurt.”, ”time”: ”2024-08-13 15:55:15”, ”Model”: ”gpt-3.5-turbo-0125”}
@@
D {”index”: 20, ”query”: ”A colony of bacteria is growing on yogurt, and the size of the colony triples every 3X days. If it takes 4Y days for the colony to completely cover the yogurt, how many days does it take for the patch to cover 1/27 of the yogurt? X and Y are both numbers, you can use them to represent the final answer. Please think step by step.”, ”answer”: ”To find out how many days it takes for the patch to cover 1/27 of the yogurt, we can first find the number of times the colony triples in size to cover the entire yogurt. If the colony triples in size every 3X days, then in 4Y days, the colony would have tripled in size 4Y / 3X times. To find out how many days it takes for the colony to cover 1/27 of the yogurt, we can use the concept of exponential growth. If the colony triples every 3X days, then after n periods of 3X days, the size of the colony would be $(3^{n})$ times the original size. So, in order to find out how many days it takes for the patch to cover 1/27 of the yogurt, we can set up the equation: $3^{(}4Y/3X)$ = 1/27 Now we need to solve for n, the number of periods of 3X days it takes for the colony to cover 1/27 of the yogurt. $3^{(}4Y/3X)=3^{(}-3)$ Since the bases are the same, we can set the exponents equal to each other. 4Y/3X = -3 Solving for Y: Y = -3X/4 So it takes -3X/4 days for the patch to cover 1/27 of the yogurt.”, ”time”: ”2024-08-14 20:51:32”, ”Model”: ”gpt-3.5-turbo-0125”}
@@
C {”index”: 20, ”query”: ”A colony of bacteria is growing on yogurt, and the size of the colony triples every three days. If it takes 27 days for the colony to completely cover the yogurt, how many days does it take for the patch to cover 1/27 of the yogurt? Please think step by step.”, ”answer”: ”Let’s think step by step through the problem: 1. We know that the size of the colony triples every three days. This means that if we represent the size of the colony at any time as ( C ), then after three days, the size of the colony will be ( 3C ). 2. Since the colony triples in size every three days, we can say that the size of the colony increases by a factor of ( 3 ) each time interval of three days passes. So, if it takes 27 days for the colony to completely cover the yogurt, then we can divide 27 by 3 to find out how many times the colony triples in size during this period: ( $frac\{27text\{days\}\}\{3text\{days/triple\}\}=9text\{triples\}$ ) This tells us that the colony triples in size 9 times over the course of 27 days. 3. Now, if the colony triples 9 times to cover the entire yogurt, we need to find out how many times it would triple to cover ( $frac\{1\}\{27\}$ ) of the yogurt. Since ( $frac\{1\}\{27\}$ ) is ( $3^{\{}-3\}$ ), or ( $frac\{1\}\{3^{3}\}$ ), we can infer that we need to triple the colony size three times fewer than the full coverage:($9text\{triples\}-3text\{triples\}=6text\{triples\}$) This means the colony needs to triple in size 6 times to cover ( $frac\{1\}\{27\}$) of the yogurt. 4. Finally, we need to calculate how many days it takes for the colony to triple in size 6 times. Since each triple takes three days, we can multiply the number of triples by three: ( 6 text{ triples} times 3 text{ days/triple} = 18 text{ days} ) So, it takes 18 days for the patch to cover ( $frac\{1\}\{27\}$) of the yogurt.”, ”time”: ”2024-08-13 15:56:43”, ”Model”: ”glm-4”}

### v2

@@ 1 Introduction
We repeated and improved upon the experiment conducted by the previous study [19], making targeted (yet as simple as possible) modifications to the original CRT problems. If LLMs genuinely possess the intrinsic capability to comprehend mathematical logic, as is widely hypothesized, their accuracy in responding to the modified problems should not experience a marked decrement. However, our experiments revealed a distinctly opposing result: even when employing the CoT approach, prominent LLMs, including the latest o1 model, continued to manifest a considerable error rate for the modified problems. Further analysis indicates that LLMs may have not developed a logic reasoning proficiency akin to System 2, nor have they acquired comprehensive mathematical cognitive abilities, including principles of mathematical operations and algorithm selection. They predominantly resort to a methodology reminiscent of System 1, which involves matching and producing responses based on the similarity between user inquiries and training data during the process of text generation. This investigation may serve to temper the overly optimistic anticipations regarding the effectiveness of CoT and the competencies of LLMs in approximating AGI.
@@ 2 Experiment I: Changing the numbers in a problem without altering its description and principle
We first replicated the experiment on the CRT3 dataset from Thilo Hagendorff’s study[19], which comprises 50 mathematical test problems. We introduced three types of modifications to the test problems, as delineated in Table 1. Type A represents the original problems in the CRT3 test, characterized as typical exponential growth problems, and incorporates three numbers in the problem statement. Type B modifies two of these numbers, while type C modifies all three. Type D replaces two numbers with letters, transforming the arithmetic problem into an algebraic one. To prevent LLMs from being unsure how to handle algebraic symbols, we appended a suffix for guidance: “X and Y are both numbers, you can use them to represent the final answer.” The modifications in Types B, C, and D do not alter the description of the original problem (Type A). Thus, the underlying mathematical principle of the original problem remains unchanged. To uncover the problem-solving strategies of LLMs, we appended explicit instructions (’Please think step by step.’) to the end of each of the four types of problems, encouraging the use of the CoT method.
@@
We evaluated five mainstream open-source and commercial models, manually analyzing and recording their accuracy in responding to the problems. A response was considered correct only if both the problem-solving approach and the final answer were accurate (refer to Appendices A ). To ensure the reliability of our experimental results, we conducted three replicates of the experiments and computed the average for each data set. The results are shown in Figure 1. In the case of Type A problems, the highest accuracy reached 100.00%, the lowest was 54.00%, and the average accuracy was 86.80%. For Type B problems, which involved modifications to two numbers, the accuracy significantly decreased, with the highest being 92.00%, the lowest 25.33%, and the average 68.53%($\delta=18.27\%;\chi^{2}(1)=23.00;P<0.001$). For the problems of Type C, where three numbers were modified, the accuracy further declined, with the highest at 80.00%, the lowest at 10.66%, and the average 53.06%($\delta=15.47\%;\chi^{2}(1)=11.91;P<0.001$). The accuracy dropped sharply once again when the arithmetic problems were transformed into algebraic ones (Type D). The best-performing generic model, GPT-4, achieved an accuracy of only 29.33%, whereas ChatGPT 3.5 scarcely provided correct responses to any problems. The other models only got around ten problems right, reducing the average accuracy down to 20.93%($\delta=32.13\%;\chi^{2}(1)=54.00;P<0.001$).
@@ 4 Experiment III: Replicating Experiment I & II on the Latest o1 Model
However, when replicating Experiment II, the average accuracy rate of o1 was only 10%, which was even lower than that of most previous LLMs. After making minimal textual modifications while altering the mathematical principles underlying the original CRT1, CRT2, and CRT3 problems, o1 persisted in selecting problem-solving approaches based on the mathematical principles corresponding to the original problems. The errors stemming from this persistence constituted 100%, 68.42%, and 75.86% of the incorrect responses to the three types of CRT problems, respectively. This may indicate that technologies such as CoT, prompt engineering, and even fine-tuning tailored to specific datasets, cannot fundamentally enhance the ability of LLMs to comprehend mathematical problems. The reason for this lies in the fact that the learning paradigm of LLMs has not undergone substantial changes (such as auto-regressive next-token predictors), leading to their thought patterns being deeply ingrained.
@@ Appendix A Examples
Problem Type Incorrect Answer Example Error Type
@@
A {”index”: 20, ”query”: ”A colony of bacteria is growing on yogurt, and the size of the colony doubles daily. If it takes 4 days for the colony to completely cover the yogurt, how many days does it take for the patch to cover half of the yogurt? Please think step by step.”, ”answer”: ”If the colony doubles in size every day, then on the first day it covers half of the yogurt. On the second day, it doubles in size and covers the other half of the yogurt, so now the yogurt is fully covered. Therefore, it takes 1 day for the colony to cover half of the yogurt.”, ”time”: ”2024-08-23 09:22:56”, ”Model”: ”gpt-3.5-turbo-0125”} calculation steps
@@
B {”index”: 20, ”query”: ”A colony of bacteria is growing on yogurt, and the size of the colony triples daily. If it takes 4 days for the colony to completely cover the yogurt, how many days does it take for the patch to cover 1/27 of the yogurt? Please think step by step.”, ”answer”: ”Let’s break down the problem step by step: 1. The colony triples daily: This means that if the initial size of the colony is 1 unit, it will be 3 units the next day, 9 units the following day, 27 units the day after, and so on. 2. It takes 4 days for the colony to completely cover the yogurt: This means that the colony reaches its maximum size on the 4th day, which is 27 units. 3. We need to find out how many days it takes for the colony to cover 1/27 of the yogurt: Since the colony triples in size daily, on the third day, the colony would be 9 units, which is 1/3 of the complete coverage. So, it means that on the third day the yogurt is 1/3 covered. Therefore, it takes 3 days for the patch to cover 1/27 of the yogurt.”, ”time”: ”2024-08-13 13:22:38”, ”Model”: ”gpt-3.5-turbo-0125”} calculation steps
@@
C {”index”: 20, ”query”: ”A colony of bacteria is growing on yogurt, and the size of the colony triples every three days. If it takes 27 days for the colony to completely cover the yogurt, how many days does it take for the patch to cover 1/27 of the yogurt? Please think step by step.”, ”answer”: ”First, let’s figure out how many times the colony triples in size in the 27 days. Since the colony triples every three days, it will triple in size 27/3 = 9 times. Now, we know that the colony starts at 1/27 of the yogurt and triples in size 9 times. So, to find out how many days it takes for the colony to cover 1/27 of the yogurt, we need to divide 27 by 9, which equals 3. Therefore, it takes 3 days for the colony to cover 1/27 of the yogurt.”, ”time”: ”2024-08-13 15:55:15”, ”Model”: ”gpt-3.5-turbo-0125”} calculation steps
@@
D {”index”: 20, ”query”: ”A colony of bacteria is growing on yogurt, and the size of the colony triples every 3X days. If it takes 4Y days for the colony to completely cover the yogurt, how many days does it take for the patch to cover 1/27 of the yogurt? X and Y are both numbers, you can use them to represent the final answer. Please think step by step.”, ”answer”: ”To find out how many days it takes for the patch to cover 1/27 of the yogurt, we can first find the number of times the colony triples in size to cover the entire yogurt. If the colony triples in size every 3X days, then in 4Y days, the colony would have tripled in size 4Y / 3X times. To find out how many days it takes for the colony to cover 1/27 of the yogurt, we can use the concept of exponential growth. If the colony triples every 3X days, then after n periods of 3X days, the size of the colony would be $(3^{n})$ times the original size. So, in order to find out how many days it takes for the patch to cover 1/27 of the yogurt, we can set up the equation: $3^{(4Y/3X)}$ = 1/27 Now we need to solve for n, the number of periods of 3X days it takes for the colony to cover 1/27 of the yogurt. $3^{(4Y/3X)}=3^{-3}$ Since the bases are the same, we can set the exponents equal to each other. 4Y/3X = -3 Solving for Y: Y = -3X/4 So it takes -3X/4 days for the patch to cover 1/27 of the yogurt.”, ”time”: ”2024-08-14 20:51:32”, ”Model”: ”gpt-3.5-turbo-0125”} calculation steps and mathematical operations
@@
[…]